1. An increase in short-run aggregate supply means ________.

A)

the real GDP would decrease and the price level would rise

B)

both the real GDP and the price level would decrease

C)

the real GDP would increase and rises in the price level would become smaller

D)

both the real GDP and rises in the price level would become greater

2. Assume that there is a 25% reserve requirement and that the Federal Reserve buys $4 billion worth of government securities. This action has the potential to increase the money supply by a maximum of ________.

A)

$14 billion

B)

$16 billion

C)

$1 billion

D)

$20 billion

3. Suppose a banking system has $ 125,000 of checkable deposits and actual reserves of $ 10,000. If the reserve ratio is 9% the banking system can expand the supply of money by a maximum of $ ____.
Answer #1

1. C) the real GDP would increase and rises in the price level would become smaller

2. B) $16 million

Increase in money supply = 4/0.25= 16

3. 98611.11

Required reserves = 0.09*12500 = 1125

Excess reserves = 10000-1125= 8875

Increase in money supply = 8875/0.09 = 98611.11